OpenStudy (anonymous):

convert root(3x) into a polar form?

OpenStudy (jdoe0001):

\(\bf y =\sqrt{3}\quad ?\)

OpenStudy (jdoe0001):

\(\bf y =\sqrt{3x}\quad ?\) rather

OpenStudy (anonymous):

y=\[\sqrt{3x}\]

OpenStudy (jdoe0001):

\(\bf y=\sqrt{3x}\\ \quad \\ \quad \\ \textit{recall that}\quad rcos(\theta)=x\qquad rsin(\theta)=y\qquad thus\\ \quad \\ y=\sqrt{3x}\implies y^2=3x\implies [rsin(\theta)]^2=3rcos(\theta)\\ \quad \\ r^2sin^2(\theta)=3rcos(\theta)\implies r=\cfrac{3\cancel{r}cos(\theta)}{\cancel{r}sin^2(\theta)}\)

OpenStudy (anonymous):

thanks